PROBLEM 2: A machine is working to a specification of 75.00±0.01 in. A study of 50 consecutive groups (of 5 pieces each) is made, yielding the following results: x 75.002 R=0.005. Find the center-line and the associated control limits (UCL and LCL) for and R control charts.
Answer #1

Since n=5 so for xbar chart we need to find ucl = xbar+A2*Rbar

Lcl = xbar-A2*Rbar

For R chart

Ucl =D4Rbar

Lcl =D3Rbar

Now for n= 5 A2= 0.577 ,D3=0 ,D4= 2.114

Using appendix table

Hence we have xbar ucl= 75.002+0.577*0.005 = 75.00485

Lcl =75.002-0.577*0.005 =74.9991

Rchart ucl = 2.114*0.005 =0.01057

Lcl =0*0.005=0